Find the zero of the polynomial : x3+ 2x2 - 5x - 6


1
Expert's answer
2021-11-29T19:40:18-0500
"x^3+ 2x^2 - 5x - 6=0"

"(x^3+x^2)+(x^2+x)+(-6x-6)=0"

"x^2(x+1)+x(x+1)-6(x+1)=0"

"(x+1)(x^2+x-6)=0"

"(x+1)(x-2)(x+3)=0"

"x_1=-3, x_2=-1, x_3=2"

Zeros: "-3, -1, 2"